OpenStudy (redneckgirl1224):

1) Work is a transfer of (1 point) energy force mass motion 2) a small 20-kilogram canoe is floation downriver at a speed of 2m/s. what is the canoes kinetic energy? (1 point) 22J 40J 80J 400J 4) which of the following is an example of an object with elastic potential energy?(1 point) a wind-up toy that has been wound up a compressed basketball a stretched rubber band all of the above 5) the energy stored in gasoline is (1 point) chemical energy electromagnetic energy mechanical energy nuclear energy

OpenStudy (redneckgirl1224):

6) what is the formula for gravitational potential energy? (1 point) KE=1/2mv^2 PE=KE PE=mgh 1.2mv^2=mgh 7) Energy which travels through space in the form of waves is known as (1point) electrical energy electromagnetic energy nuclear energy thermal ennergy

OpenStudy (redneckgirl1224):

Will give a medal for best answer!!!!!!

OpenStudy (fortisdea13):

1. A When a force is applied to an object while the object moves with some aligned component, work is done on the object. 2. B, 40J Ek=1/2m(v)^2 <-- formula Ek= Kinetic Energy! Ek=1/2(20kg)( 2m/s)^2. Ek=1/2(20kg)(4m^2/s2) = .5*20gk*4m^2/s^2 4. D, All of the above

OpenStudy (fortisdea13):

5. A, Chemical energy 6. C, PE=mgh 7. B, Electromagnetic energy

OpenStudy (redneckgirl1224):

THANK YOU!!!!!!